OpenStudy (anonymous):

(9x over 7)-x=x-28 over 21 solve the equation

OpenStudy (anonymous):

\[(\frac{ 9x }{ 7 }) - x = (\frac{ x-28 }{ 21 })\] First, find the least common denominator, which is 21. Then multiply that through for each term:\[(21)(\frac{ 9x }{ 7 }) - (21)x = (21) (\frac{ x-28 }{ 21 })\] 27x - 21x = x - 28 6x = x - 28 5x = -28 x = -28/5

OpenStudy (anonymous):

thANK YOU